I need help with the following problems:

1) When aqueous solutions of potassium carbonate and aluminum chloride are combined, solid aluminum carbonate and a solution of potassium chloride are formed. The net ionic equation for this reaction is:

2 )When aqueous solutions of sodium nitrite and hydrochloric acid are mixed, an aqueous solution of sodium chloride and nitrous acid results. Write the net ionic equation for the reaction.

3) Does a reaction occur when aqueous solutions of sodium carbonate and ammonium sulfide are combined? If a reaction does occur, write the net ionic equation.

4) Does a reaction occur when aqueous solutions of ammonium carbonate and sodium iodide are combined? If a reaction does occur, write the net ionic equation.

5) Does a reaction occur when aqueous solutions of barium hydroxide and magnesium chloride are combined? If a reaction does occur, write the net ionic equation.

6) Does a reaction occur when aqueous solutions of nickel(II) nitrate and ammonium carbonate are combined? If a reaction does occur, write the net ionic equation.

7) Write a net ionic equation for the reaction that occurs when aqueous solutions of perchloric acid and sodium hydroxide are combined. (Use H+ instead of H3O+.)

8) Write a net ionic equation for the reaction that occurs when aqueous solutions of nitrous acid and potassium hydroxide are combined.
Answer #1

Solution:

1) Potassium carbonate + Aluminium chloride

Molecular equation will be:

3K2CO3(aq) + 2AlCl3 (aq) = 6KCl (aq) + Al2 (CO3)3 (S)

Net ionic equation will be:

3CO3^2- (aq) + 2Al^3+ (aq) = Al2(CO3)3 (s)

2) Sodium nitrite + Hydrochloric acid

NaNO2(aq) + HCl (aq) = NaCl (aq) + HNO2(aq)

Net ionic equation will be:

NO2^- (aq) + H+(aq) = HNO2(aq)

3) Ammonium sulfide does not react with sodium carbonate because both behave as base.

4) Ammonium carbonate and sodium iodide do not react with each other because both are salts.